OpenStudy (anonymous):

solve for 0<θ<2(3.14[pi]): 9cos^((2))θ-2cosθ-3=0 give answer in terms of 3.14[pi]

sam (.sam.):

Since it can't be factored, you need to use the quadratic formula. \[\cos(\theta)=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\] \[\cos(\theta)=\frac{-(-2) \pm \sqrt{(-2)^2-4(9)(-3)}}{2(9)}\] \[\cos(\theta)=\frac{2 \pm \sqrt{112}}{18}\] ------------------------------------------- \[\cos(\theta)=\frac{2 + \sqrt{112}}{18}~~~\text{and}~~~\cos(\theta)=\frac{2 - \sqrt{112}}{18}\] ----------------------------------------- Take the inverse of both trigo and get \(\theta\)

sam (.sam.):

You'll need this too